OpenStudy (maheshmeghwal9):

Solve:

OpenStudy (maheshmeghwal9):

\[x \left| x-4 \right|+a=0\]

OpenStudy (anonymous):

what have you tried so far?

OpenStudy (maheshmeghwal9):

i have done \[x-4=\pm (-a/x)\]

OpenStudy (anonymous):

2-sqrt(4-a), sqrt(4-a)+2, 2-sqrt(a+4), sqrt(a+4)+2

OpenStudy (maheshmeghwal9):

How?

OpenStudy (maheshmeghwal9):

please tell

OpenStudy (anonymous):

it is quadratic

OpenStudy (maheshmeghwal9):

k!

OpenStudy (anonymous):

(-b±sqrt(b^2-4ac))/2a

OpenStudy (anonymous):

no wait it can be done much easier

OpenStudy (maheshmeghwal9):

But please tell how can I make original eq. into quadratic

OpenStudy (anonymous):

i think it can be \[(x-4)^2x^2=(-a)^2\]

OpenStudy (maheshmeghwal9):

k!

OpenStudy (anonymous):

ok x−4=±(−a/x) x(x-4) = ±(−a) x^2-4x = ±(−a) x^2-4x±a = 0 using the formula x = (4±sqrt(16±4a))/2 x = (4±sqrt(4(4±a)))/2 x = (4±2sqrt(4±a))/2 x = 2±sqrt(4±a)

OpenStudy (anonymous):

there are 4 possible solutions

OpenStudy (maheshmeghwal9):

oh thanx a lot! I got it now.